• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name johna

    Log entry time 16:03:41 on October23,2002

    Entry number 88886

    keyword=Shift summary

    Shift summary, day shift, Wednesday October 23rd

    8:00 Beam stability problems from end of last shift disappear at 8:00 on the nose. Stable 100 uA beam at start of shift.

    ~9:55 Lost communication with the target IOC. Target temperature was stable until the beam went away at ~10:00. This was during run 2702, but since the fluctuations did not occur until the beam went away, the data should be OK . The IOC rebooted itself and the target is back to normal.

    10:00 MCC takes beam, we request controlled access.

    ~11:45 Work is finished in the hall (target problems fixed), ask for beam permit and beam

    12:20 Back at 110 uA, start run 2703

    ~12:40 Asked for beam to be centered in x (during run 2705)

    ~1:00 During shift checklist, noticed that FFB was off. MCC says that it was turned off at 8am this morning (which is what fixed the trip/stability problems they were having). They are waiting on additional diagnostics before turning it back on.

    ~13:24 MCC called to tell us that they have turned on the energy lock and

    14:10 Gas alarm - out of nitrogen. Jack is changing the bottles, estimates
    that the VDCs won't notice, the FPPs might notice the brief interruption. Occurred near the end of run 2706.

    15:15 Finish run 2707. Ask for beam off for 5 minutes so that Chris Keith can adjust the high powered heater settings.

    15:25 Starting run 2708.

    15:35 Q1 trips. Ed is going down to reset it.

    Runs taken: All are Q3_d40 kinematics

    2699: 1M events, nice stable beam (after first 15 min). I_ave=88 uA
    2700: 1M events, stable beam, I_ave=94
    2701: Scintallator efficiency run (ps1-ps5=1), 1M events
    2702: 50k events, stable beam (ended at 10:00 when MCC took beam away)
    2703&2704: Junk and/or Cosmics
    2705: 1M events, fairly stable beam (position drifting at the 0.2-0.3mm level).
    2706: 1M events, stable beam (FFB back on at ~100k events).
    2707: 1M events, stable beam, FFB on, beam centered.
    2708: 56k events, ends with Q1 trip of left arm.